OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 2y = -6 3x - 2y = 11 What is the solution to the system? (-1, -7) (1, -4) (2, -1) (3, -2)

OpenStudy (anonymous):

x = - 3 - y then you substitute x on the second equation, and you should be able to solve it. Or you can just add both equations. either way its ok

OpenStudy (anonymous):

first solve in terms of x or y. you can pick any equation. 2x+2y=-6 2x=-6-2y x=-3-y then sub (-3-y) for x in second equation to solve for y. once you get y the rest should be straightforward.

hero (hero):

2x + 2y = -6 3x - 2y = 11 Hint: Add the two equations together 5x = 5

hero (hero):

x = 5/5 x = 1

hero (hero):

Only one solution where x = 1

hero (hero):

But you can solve for y to make sure. Another way to check is to substitute the (x,y) point you believe is correct into the original equations